# Proximity-induced superconductivity and emerging topological phases in altermagnet-based heterostructures

### v1

@@ Abstract
We present a theoretical framework for investigating superconducting proximity effect in altermagnet (AM)–superconductor (SC) heterostructures. In general, AMs, characterized by vanishing net magnetization but spin-split electronic spectra, provide a promising platform for realizing unconventional magnetic phases. We consider a two-dimensional $d$-wave AM proximity coupled to a three dimensional ordinary $s$-wave SC. By integrating out the superconducting degrees of freedom, we derive an effective Hamiltonian that describes the proximity-induced modifications in the AM layer in the form of a self-energy. We then derive an effective Green’s function to obtain the proximity-induced pairng amplitudes in the AM layer and classify the induced pairing amplitudes according to their parity, frequency, and spin. We find the presence of even-parity singlet and triplet pairing amplitudes in the AM layer. To achieve the odd-parity triplet components, important to realize topological superconductivity, we introduce a layer of Rashba spin-orbit coupling (RSOC) in the heterostructure. We analyse the band topology of this proximity-induced AM-RSOC layer and demonstrate the emergence of both weak and strong topological superconducting phases with edge-localized modes, characterized by winding number and Chern number. These findings highlight the role of AM–SC hybrid setup as a versatile platform for realizing odd-parity triplet pairings and engineering topological superconductivity in two-dimension.
@@ I Introduction
The recent discovery of a novel class of magnetic materials, termed as altermagnets (AMs), lead to a plethora of research activity over the past few years  Šmejkal et al. (2022a, b); Mazin (2022); Bhowal and Spaldin (2024); Mazin (2023). These AMs feature collinear compensated magnetic order where the opposite spin sublattices are related via rotation rather than translation or inversion. Unlike the conventional antiferromagnets, the broken time-reversal symmetry (TRS) in AMs leads to momentum dependent spin-splitting of the electronic bands inspite of vanishing net magnetization. In two dimensions (2D), the most commonly used AM is the $d$-wave AM where the magnetic order mimics the shape of planar $d$-wave orbitals with the up and down spin sectors are connected via $C_{4}$ rotation about the axis perpendicular to the plane. Candidate materials for such AMs include RuO2 Šmejkal et al. (2022a, b); Zhou et al. (2024); Mazin (2022); Bai et al. (2023), MnF2 Bhowal and Spaldin (2024), and MnTe Lee et al. (2024); Mazin (2023) etc. Subsequently, substantial development have been proposed in unconventional odd-parity noncollinear magnetic systems, termed as $p$-wave magnets Brekke et al. (2024); Nagae et al. (2025); Hellenes et al. ; Maeda et al. (2024a). However, in contrast to AMs, $p$-wave magnets preserve TRS but break parity symmetry leading to spin-splitted band structure.
@@
Over the past few decades, superconductvity has always been a major area of research activity for realizing intricate phenomena. One of the fascinating phenomena is the superconducting proximity effect in mesoscopic systems where a non-vanishing supeconducting order is induced in a non-superconducitng material placed in close proximity to a bulk superconductor (SC) Deutscher and de Gennes (1969); Bergeret et al. (2005); Buzdin (2005). Microscopically, the superconducting proximity effect emerges due to phase coherence between electron and hole wavefunctions even outside the superconducting region Beenakker (1997) leading to tunnelling of Cooper pairs into a nonsuperconducting material, thereby inducing superconducting correlations over some characteristic length scales. Proximity effects not only enable superconductivity in otherwise nonsuperconducting systems, but also strongly modify superconducting properties near interfaces. This gives rise to rich physics in a wide range of hybrid junctions-where the induced Cooper-pair symmetries have been extensively analyzed in the literature Buzdin (2005); Fu and Kane (2008); Burset et al. (2015); Chevallier et al. (2013); Khanna et al. (2014); Sitthison and Stanescu (2014); Linder et al. (2010); Black-Schaffer and Balatsky (2013a, b); Dutta and Black-Schaffer (2019); Dutta et al. (2020).
@@
The detrimental interplay between magnetism and superconductivity has long been recognized as a fertile platform for realizing unconventional quantum phases. In this direction, focussing on AM-SC heterostructures, several theoretical investigation have been carried out which includes Andreev reflections Maeda et al. (2024b); Papaj (2023); Sun et al. (2023); Zhao et al. (2025), Josephson effects Ouassou et al. (2023); Beenakker and Vakhtel (2023); Cheng and Sun (2024); Pal et al. ; Fukaya et al. (2025); Cheng et al. (2024); Chakraborty and Black-Schaffer (a); Debnath et al. , magnetoelectric responses Zyuzin ; Hu et al. (2025), finite-momentum pairing Zhang et al. (2024); Chakraborty and Black-Schaffer (b), and routes to engineer topological superconductivity Ghorashi et al. (2024); Mondal et al. (2025); Li et al. (2024); Li and Liu (2023); Zhu et al. (2023); Li (2024). A central goal in this context is the realization of topological suprconductor (TSC), hosting Majorana modes as zero-energy quasiparticle excitations Kitaev (2001); Qi and Zhang (2011); Leijnse and Flensberg (2012); Alicea (2012); Nayak et al. (2008). These neutral excitations are central to the proposals for fault-tolerant topological quantum computation owing to their non-Abelian braiding statistics Ivanov (2001); Freedman et al. (2003); Stern (2010). While numerous engineered platforms have been proposed in one-dimension (1D) and 22Lutchyn et al. (2010); Das et al. (2012); Aguado (2017); Nadj-Perge et al. (2013); Yazdani (2015); Pientka et al. (2013); Mondal et al. (2023a); Mourik et al. (2012); Beenakker (2013); Fu and Kane (2008); Chatterjee et al. (2024)–mostly rely on magnetic impurities or external magnetic fields, which tend to suppress superconductivity. On the other hand, AMs offer an appealing alternative in 22D as they provide a source of broken TRS prevailing larger bulk superconducting gaps compared to the other platforms to realize TSC.
@@
Despite these advantages, most theoretical works so far treated the induced superconductivity phenomenologically Ghorashi et al. (2024); Mondal et al. (2025); Maeda et al. (2025); Li and Liu (2023), by inserting a constant pairing term directly into the low-energy Hamiltonian of the non-superconducting system. While such models capture qualitative aspects, they neglect the microscopic details of tunneling and interface processes that are crucial for determining the induced pairing structure and its robustness for the realization of TSC and related phenomena.
In this work, we present a microscopic theory of superconducting proximity effects in AM–SC heterostructures. Specifically, we investigate a system consisting of a 22D $d$-wave AM placed on top of a three-dimensional (3D) conventional $s$-wave SC. By integrating out the superconducting degrees of freedom Chevallier et al. (2013); Khanna et al. (2014), we derive an effective Hamiltonian and Green’s function therein for the AM layer that incorporates tunneling processes explicitly. This framework allows us to: (i) perform spectral analysis and (ii) characterize induced pairing symmetries, including mixed even- and odd-frequency channels. Additionally, we investigate the role of Rashba spin–orbit coupling (RSOC) in generating odd-parity triplet components, and analyse the emergence of weak and strong topological superconductivity in this hybrid setup. Using exact diagonalization and Chern number calculation Fukui et al. (2005), we demonstrate that the AM–SC platform supports boundary Majorana edge modes (MEMs) characterized by appropriate topological invariants. Importantly, the tunneling amplitude appears as an experimentally tunable parameter that directly connects our microscopic description to realistic devices.
@@ II Model and Method
$H_{AM}$ $=\mu\sum_{\mathbf{r}}\Psi_{\mathbf{r}}^{\dagger}(\tau_{z}\sigma_{0})\Psi_{\mathbf{r}}-t\sum_{\langle\mathbf{r},\mathbf{r}^{\prime}\rangle}\Psi_{\mathbf{r}}^{\dagger}(\tau_{z}\sigma_{0})\Psi_{\mathbf{r}^{\prime}}$ (1)
@@
where, $\Psi_{\mathbf{r}}=\begin{pmatrix}\psi_{\uparrow},\psi_{\downarrow},\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger}\end{pmatrix}_{\mathbf{r}}^{T}$ is the Nambu spinor in the BdG basis with $\psi_{\mathbf{r}s}~(\psi_{\mathbf{r}s}^{\dagger})$ correponds to the electron annhilation (creation) operator at the site $\mathbf{r}=(x,y)$ with spin $s=(\uparrow,\downarrow)$ in the altermagnetic layer. Here, $\sigma$ and $\tau$ represent Pauli matrices in spin and particle-hole space, respectively and $\mu$, $t$ and $J_{a}$ denote the chemical potential, nearest-neighbor hopping amplitudes, and AM exchange energy respectively.
@@
$H_{SC}$ $=\mu_{\text{sc}}\sum_{\mathbf{R},\sigma}\Phi_{\mathbf{R},\sigma}^{\dagger}\Phi_{\mathbf{R},\sigma}-t_{\text{sc}}\sum_{\langle\mathbf{R},\mathbf{R}^{\prime}\rangle,\sigma}\Phi_{\mathbf{R},\sigma}^{\dagger}\Phi_{\mathbf{R}^{\prime},\sigma}$ (2)
@@
Since the Hamiltonian in the SC (Eq. (2)) is quadratic in the fermionic degrees of freedom, we integrate it out to obtain an effective action for the AM. Following the procedure for computing the self-energy as outlined in Refs.Khanna et al. (2014); Chevallier et al. (2013); Cuevas et al. (1996), we decouple the AM and SC degrees of freedom and calculate the Green’s function $G(\omega)$ for the altermagnetic layer (a sketch of the derivation is provided in Appendix A) as,
@@
Here, $\lambda_{s}=\pi N(0)|\tilde{t}|^{2}$ characterizes the effective coupling strength to the SC, with $N(0)$ denoting the normal-state DOS at the Fermi level of the SC. In Eq. (6), $\eta$ refers to a positive infinitesimal number that ensures the causality of the Green’s function.
@@ III DOS and Pairing Amplitude
The anomalous Green’s function encodes the SC correlations in the AM system. To identify the nature of these correlations, it is necessary to analyze the symmetry properties of the pair amplitude $\Delta_{\sigma\sigma^{\prime}}(\mathbf{k},\omega)$. This classification is guided by the antisymmetry constraint imposed by the Fermi–Dirac statistics on the pair amplitude under the total exchange of the quantum numbers Tanaka et al. (2012); Cayao et al. (2020); Maeda et al. (2025); Dutta et al. (2021) as
@@
In order to carry out the calculation of frequency-resolved pairing amplitudes, we use the retarded Green’s functions for $\omega<0$ and the advanced Green’s functions for $\omega>0$ Parhizgar and Black-Schaffer (2020). By inspection of Eq. (14), the spin-singlet component $\psi(\mathbf{k},\omega)$ is an even function of both frequency and momentum. This follows from the fact that $\chi(\omega,\mathbf{k})=\chi(\omega,-\mathbf{k})$, and is further corroborated by the results presented in Fig. 3(a), where both the real and imaginary parts of $\psi(\mathbf{k},\omega)$ remain even as a function of $\omega$. We therefore identify $\psi$ as belonging to the ESE class. In contrast, from Eq. (14), it is clear that the momentum dependence of the spin-triplet component $d_{z}(\mathbf{k},\omega)$ arises from $\chi(\omega,\mathbf{k})$ and $h_{\mathrm{AM}}(\mathbf{k})$. Note that, both of them are even functions of $\mathbf{k}$. As shown in Fig. 3(b), both the real and imaginary parts of $d_{z}(\mathbf{k},\omega)$ are odd functions of $\omega$. Hence, $d_{z}(\mathbf{k},\omega)$ belongs to the OTE class, consistent with the antisymmetry condition of Eq. (10). Importantly, $d_{z}(\mathbf{k},\omega)$ originates from the $d$-wave altermagnetic field $h_{\mathrm{AM}}(\mathbf{k})$, which acts as a spin-mixing mechanism enabling triplet correlations in the AM layer (a similar pairing symmetry analysis for the newly proposed $p$-wave magnet is presented in Appendix D). To illustrate these features, Fig. 3(c) displays the absolute value of the frequency resolved OTE amplitude $|d_{z}|$. Four nodal lines appear along the diagonals, reflecting the underlying $d$-wave structure of the AM. Thus, while ESE pairing is directly inherited from the parent $s$-wave SC, OTE pairing appears from the interplay between the $d$-wave AM and the spin-singlet SC. The relative weight of the induced triplet component is quantified in Fig. 3(d), which manifests the ratio of the integrated squared magnitudes of OTE and ESE amplitudes across the $(J_{a},\omega)$ plane. A finite ratio over a broad parameter range demonstrates the coexistence of both the components. Notably, OTE pairing becomes dominant when $J_{a}\sim\lambda_{s}/2$, as indicated by the red regions. Although symmetry analysis categorizes $d_{z}$ as OTE, i.e., even in parity, a finite $p$-wave–like structure usually requires odd parity pairing. In the following, we explore the role of RSOC in the AM layer as a natural mechanism that can promote such momentum-dependent triplet terms.
@@
Figure 4: Momentum-resolved pairing amplitudes are depicted in the $k_{x}$-$k_{y}$ plane choosing $\omega=0.5\Delta$. Panels (a) and (b) show the real and imaginary parts of the singlet component $\psi(\mathbf{k})$. On the other hand, panels (c)–(h) display the real and imaginary parts of the triplet components $d_{x}(\mathbf{k})$, $d_{y}(\mathbf{k})$, and $d_{z}(\mathbf{k})$. The finite in-plane triplet components $d_{x}$, $d_{y}$ arise in the presence of RSOC $\alpha=0.3\Delta$]. The other model parameters used are $J_{a}=0.1\Delta$, $\mu=0t$, $\Delta=0.3t$, $\lambda_{s}=0.5\Delta$.

### v2

@@ Abstract
We present a theoretical framework for investigating superconducting proximity effect in altermagnet (AM)–superconductor (SC) heterostructures. In general, AMs, characterized by vanishing net magnetization but spin-split electronic spectra, provide a promising platform for realizing unconventional quantum phases. We consider a two-dimensional $d$-wave AM proximity coupled to a three dimensional ordinary $s$-wave SC. By integrating out the superconducting degrees of freedom, we derive an effective Hamiltonian that describes the proximity-induced modifications in the AM layer in the form of a self-energy. We then derive an effective Green’s function to obtain the proximity-induced pairing amplitudes in the AM layer and classify the induced pairing amplitudes according to their parity, frequency, and spin. We find the presence of even-parity singlet and triplet pairing amplitudes in the AM layer. To achieve the odd-parity triplet components, important to realize topological superconductivity, we introduce a layer of Rashba spin-orbit coupling (RSOC) in the heterostructure. We analyse the band topology of this proximity-induced AM-RSOC layer and demonstrate the emergence of both weak and strong topological superconducting phases with edge-localized modes, characterized by winding number and Chern number. These findings highlight the role of AM–SC hybrid setup as a versatile platform for realizing odd-parity triplet pairings and engineering topological superconductivity in two dimensions.
@@ I Introduction
The recent discovery of a novel class of magnetic materials, termed as altermagnets (AMs), lead to a plethora of research activity over the past few years  Šmejkal et al. (2022a, b); Mazin (2022); Bhowal and Spaldin (2024); Mazin (2023). These AMs feature collinear compensated magnetic order where the opposite spin sublattices are related via rotation rather than translation or inversion. Unlike the conventional antiferromagnets, the broken time-reversal symmetry (TRS) in AMs leads to momentum dependent spin-splitting of the electronic bands inspite of vanishing net magnetization. In two dimensions (2D), the most commonly used AM is the $d$-wave AM where the magnetic order mimics the shape of planar $d$-wave orbitals with the up and down spin sectors are connected via $C_{4}$ rotation about the axis perpendicular to the plane. Candidate materials for such AMs include RuO2 Šmejkal et al. (2022a, b); Zhou et al. (2024); Mazin (2022); Bai et al. (2023), MnF2 Bhowal and Spaldin (2024), and MnTe Lee et al. (2024); Mazin (2023) etc. Subsequently, substantial development have been proposed in unconventional odd-parity noncollinear magnetic systems, termed as $p$-wave magnets Brekke et al. (2024); Nagae et al. (2025); Hellenes et al. ; Maeda et al. (2024); Fukaya et al. (2025a). However, in contrast to AMs, $p$-wave magnets preserve TRS but break parity symmetry leading to spin-splitted band structure.
@@
Over the past few decades, superconductvity has always been a major area of research activity for realizing intricate phenomena. One of the fascinating phenomena is the superconducting proximity effect in mesoscopic systems where a non-vanishing supeconducting order is induced in a non-superconducitng material placed in close proximity to a bulk superconductor (SC) Deutscher and de Gennes (1969); Bergeret et al. (2005); Buzdin (2005). Microscopically, the superconducting proximity effect emerges due to phase coherence between electron and hole wavefunctions even outside the superconducting region Beenakker (1997) leading to tunnelling of Cooper pairs into a nonsuperconducting material, thereby inducing superconducting correlations over some characteristic length scales. Proximity effects not only enable superconductivity in otherwise nonsuperconducting systems, but also strongly modify superconducting properties near interfaces. This gives rise to rich physics in a wide range of hybrid junctions-where the induced Cooper-pair symmetries have been extensively analyzed in the literature Buzdin (2005); Fu and Kane (2008); Burset et al. (2015); Chevallier et al. (2013); Khanna et al. (2014); Sitthison and Stanescu (2014); Linder et al. (2010); Black-Schaffer and Balatsky (2013a, b); Dutta and Black-Schaffer (2019); Dutta et al. (2020); Eschrig (2015).
@@
The detrimental interplay between magnetism and superconductivity has long been recognized as a fertile platform for realizing unconventional quantum phases. In this direction, focussing on AM-SC heterostructures Fukaya et al. (2025b), several theoretical investigation have been carried out which includes Andreev reflections Papaj (2023); Sun et al. (2023); Zhao et al. (2025), Josephson effects Ouassou et al. (2023); Beenakker and Vakhtel (2023); Cheng and Sun (2024); Pal et al. (a); Fukaya et al. (2025c); Cheng et al. (2024); Chakraborty and Black-Schaffer (2025a); Debnath et al. , magnetoelectric responses Zyuzin ; Hu et al. (2025), finite-momentum pairing Zhang et al. (2024); Chakraborty and Black-Schaffer (2024); Sumita et al. (2025); Chakraborty and Black-Schaffer (2025b); Rasmussen et al. (2025), and routes to engineer topological superconductivity Ghorashi et al. (2024); Mondal et al. (2025); Li et al. (2024); Li and Liu (2023); Zhu et al. (2023); Li (2024); Pal et al. (b). A central goal in this context is the realization of topological suprconductor (TSC), hosting Majorana modes as zero-energy quasiparticle excitations Kitaev (2001); Qi and Zhang (2011); Leijnse and Flensberg (2012); Alicea (2012); Nayak et al. (2008). These neutral excitations are central to the proposals for fault-tolerant topological quantum computation owing to their non-Abelian braiding statistics Ivanov (2001); Freedman et al. (2003); Stern (2010). While numerous engineered platforms have been proposed in one-dimension (1D) and 22Lutchyn et al. (2010); Das et al. (2012); Aguado (2017); Nadj-Perge et al. (2013); Yazdani (2015); Pientka et al. (2013); Mondal et al. (2023a); Mourik et al. (2012); Beenakker (2013); Fu and Kane (2008); Chatterjee et al. (2024)–mostly rely on magnetic impurities or external magnetic fields, which tend to suppress superconductivity. On the other hand, AMs offer an appealing alternative in 22D as they provide a source of broken TRS prevailing larger bulk superconducting gaps compared to the other platforms to realize TSC.
@@
Despite these advantages, most theoretical works so far treated the induced superconductivity phenomenologically Ghorashi et al. (2024); Mondal et al. (2025); Maeda et al. (2025); Li and Liu (2023); Fukaya et al. (2025b), by inserting a constant pairing term directly into the low-energy Hamiltonian of the non-superconducting system. While such models capture qualitative aspects, they neglect the microscopic details of tunneling and interface processes that are crucial for determining the induced pairing structure and its robustness for the realization of TSC and related phenomena.
In this work, we present a microscopic theory of superconducting proximity effects in AM–SC heterostructures. Specifically, we investigate a system consisting of a 22D $d$-wave AM placed on top of a three-dimensional (3D) conventional $s$-wave SC. By integrating out the superconducting degrees of freedom Chevallier et al. (2013); Khanna et al. (2014), we derive an effective Hamiltonian and Green’s function therein for the AM layer that incorporates tunneling processes explicitly. This framework allows us to: (i) perform spectral analysis and (ii) characterize induced pairing symmetries, including mixed even- and odd-frequency channels. Additionally, we investigate the role of interfacial Rashba spin–orbit coupling (RSOC) in generating odd-parity triplet components Shen et al. (2024), and analyse the emergence of weak and strong topological superconductivity in this hybrid setup. Using exact diagonalization and Chern number calculation Fukui et al. (2005), we demonstrate that the AM–SC platform supports boundary Majorana edge modes (MEMs) characterized by appropriate topological invariants. Importantly, the tunneling amplitude appears as an experimentally tunable parameter that directly connects our microscopic description to realistic devices.
@@ II Model and Method
$H_{AM}$ $=-\mu\sum_{\mathbf{r}}\Psi_{\mathbf{r}}^{\dagger}(\tau_{z}\sigma_{0})\Psi_{\mathbf{r}}-t\sum_{\langle\mathbf{r},\mathbf{r}^{\prime}\rangle}\Psi_{\mathbf{r}}^{\dagger}(\tau_{z}\sigma_{0})\Psi_{\mathbf{r}^{\prime}}$ (1)
@@
where, $\Psi_{\mathbf{r}}=\begin{pmatrix}\psi_{\uparrow},\psi_{\downarrow},\psi_{\downarrow}^{\dagger},-\psi_{\uparrow}^{\dagger}\end{pmatrix}_{\mathbf{r}}^{T}$ is the Nambu spinor in the BdG basis with $\psi_{\mathbf{r}s}~(\psi_{\mathbf{r}s}^{\dagger})$ correponds to the electron annhilation (creation) operator at the site $\mathbf{r}=(x,y)$ with spin $s=(\uparrow,\downarrow)$ (actually pseudospin in case of AM  Chakraborty and Black-Schaffer (2025b)) in the altermagnetic layer. Here, $\sigma$ and $\tau$ represent Pauli matrices in spin and particle-hole space, respectively and $\mu$, $t$ and $J_{a}$ denote the chemical potential, nearest-neighbor hopping amplitudes, and AM exchange energy respectively.
@@
$H_{SC}$ $=-\mu_{\text{sc}}\sum_{\mathbf{R},\sigma}\Phi_{\mathbf{R},\sigma}^{\dagger}\Phi_{\mathbf{R},\sigma}-t_{\text{sc}}\sum_{\langle\mathbf{R},\mathbf{R}^{\prime}\rangle,\sigma}\Phi_{\mathbf{R},\sigma}^{\dagger}\Phi_{\mathbf{R}^{\prime},\sigma}$ (2)
@@
Since the Hamiltonian in the SC (see Eq. (2)) is quadratic in the fermionic degrees of freedom, we integrate it out to obtain an effective action for the AM. This allows us to explicitly focus on the altermagnetic layer as the physical subsystem of interest, with the superconducting degrees of freedom participating only through a self-energy term that encodes proximity-induced superconducting correlations in the AM. Following the procedure for computing the self-energy as outlined in Refs. Khanna et al. (2014); Chevallier et al. (2013); Cuevas et al. (1996), we decouple the AM and SC degrees of freedom and calculate the Green’s function $G(\omega)$ for the altermagnetic layer (a sketch of the derivation is provided in Appendix A) as,
@@
Here, $\lambda_{s}=\pi N(0)|\tilde{t}|^{2}$ characterizes the effective coupling strength to the SC, with $N(0)$ denoting the normal-state DOS at the Fermi level of the SC. In Eq. (6), $\eta$ refers to a positive infinitesimal number that ensures the causality of the Green’s function. The coupling strength between the AM and SC layer is primarily controlled by the $N(0)$, and $\tilde{t}$. Note that, $N(0)$ depends on the model parameters of the SC, such as hopping amplitude and Fermi energy of the SC, thus cannot take arbitrary values. In addition, the strength of $\tilde{t}$ depends on the interfacial properties of the AM-SC layer which we assume to be weak compared to other model parameters $t$, $\Delta$. For this reasons, the strength of $\lambda_{s}$ is constrained by $N(0)$ and $\tilde{t}$, and cannot take arbitrary large values within the weak-coupling regime. Additionally, it seems that for $\lambda_{s}>\Delta$, the proximity-induced gap becomes larger than the superconducting gap of the parent SC. However, physically such enhancement is not possible as our formalism is mainly valid in the weak coupling regime i.e., $\lambda_{s}<\Delta$. Thus, in particular, one must avoid parameter values for which $\lambda_{s}$ becomes larger than $\Delta$. To ensure that, all numerical results presented in our manuscript are restricted to the regime $\lambda_{s}\leq 0.5\Delta$.
@@ III DOS and Pairing Amplitude
The anomalous Green’s function encodes the SC correlations in the AM system. To identify the nature of these correlations, it is necessary to analyze the symmetry properties of the pair amplitude $\Delta_{\sigma\sigma^{\prime}}(\mathbf{k},\omega)$. This classification is guided by the antisymmetry constraint imposed by the Fermi–Dirac statistics on the pair amplitude under the total exchange of the quantum numbers Tanaka et al. (2012); Cayao et al. (2020); Maeda et al. (2025); Dutta et al. (2021); Fukaya et al. (2025b) as
@@
In order to carry out the calculation of frequency-resolved pairing amplitudes, we use the retarded Green’s functions for $\omega<0$ and the advanced Green’s functions for $\omega>0$ Parhizgar and Black-Schaffer (2020). By inspection of Eq. (14), the spin-singlet component $\psi(\mathbf{k},\omega)$ is an even function of both frequency and momentum. This follows from the fact that $\chi(\omega,\mathbf{k})=\chi(\omega,-\mathbf{k})$, and is further corroborated by the results presented in Fig. 3(a), where both the real and imaginary parts of $\psi(\mathbf{k},\omega)$ remain even as a function of $\omega$. We therefore identify $\psi$ as belonging to the ESE class. In contrast, from Eq. (14), it is clear that the momentum dependence of the spin-triplet component $d_{z}(\mathbf{k},\omega)$ arises from $\chi(\omega,\mathbf{k})$ and $h_{\mathrm{AM}}(\mathbf{k})$. Note that, both of them are even functions of $\mathbf{k}$. As shown in Fig. 3(b), both the real and imaginary parts of $d_{z}(\mathbf{k},\omega)$ are odd functions of $\omega$. Hence, $d_{z}(\mathbf{k},\omega)$ belongs to the OTE class, consistent with the antisymmetry condition of Eq. (10). Importantly, $d_{z}(\mathbf{k},\omega)$ originates from the $d$-wave altermagnetic field $h_{\mathrm{AM}}(\mathbf{k})$, which acts as a spin-mixing mechanism enabling triplet correlations in the AM layer (a similar pairing symmetry analysis for the newly proposed $p$-wave magnet is presented in Appendix D). To illustrate these features, Fig. 3(c) displays the absolute value of the frequency resolved OTE amplitude $|d_{z}|$. Four nodal lines appear along the diagonals, reflecting the underlying $d$-wave structure of the AM. Such nodal-line structure originates from the characteristic momentum-dependent spin splitting of AMs and constitutes a qualitative distinction from conventional ferromagnetic systems, where such nodal features are generically absent. Thus, while ESE pairing is directly inherited from the parent $s$-wave SC, OTE pairing appears from the interplay between the $d$-wave AM and the spin-singlet SC. The relative weight of the induced triplet component is quantified in Fig. 3(d), which manifests the ratio of the integrated squared magnitudes of OTE and ESE amplitudes across the $(J_{a},\omega)$ plane. A finite ratio over a broad parameter range demonstrates the coexistence of both the components. Notably, OTE pairing becomes dominant when $J_{a}\sim\lambda_{s}/2$, as indicated by the red regions. Although symmetry analysis categorizes $d_{z}$ as OTE, i.e., even in parity, a finite $p$-wave–like structure usually requires odd parity pairing. In the following, we explore the role of RSOC in the AM layer as a natural mechanism that can promote such momentum-dependent triplet terms.
@@
Figure 4: Momentum-resolved pairing amplitudes are depicted in the $k_{x}$-$k_{y}$ plane choosing $\omega=0.5\Delta$. Panels (a) and (b) show the real and imaginary parts of the singlet component $\psi(\mathbf{k})$. On the other hand, panels (c)–(h) display the real and imaginary parts of the triplet components $d_{x}(\mathbf{k})$, $d_{y}(\mathbf{k})$, and $d_{z}(\mathbf{k})$. The finite in-plane triplet components $d_{x}$, $d_{y}$ arise in the presence of RSOC $\alpha=0.3\Delta$]. The other model parameters used are $J_{a}=0.1\Delta$, $\mu=0$, $\Delta=0.3t$, $\lambda_{s}=0.5\Delta$.